OpenStudy (anonymous):

Updated Batting Averages are computed using the formula: A = B+K/G A = batting average B = times at bat multiplied by player’s seasonal average K = hits in the new game G = times at bat plus times at bat in a new game Solve A = for K. Please show your work. (6 points) Use your new formula and the values given in the table below to find the player’s hits in a new game (K). Round your answer to the near

OpenStudy (anonymous):

@best.shakir Can You Help Me?

OpenStudy (anonymous):

need the values from the table you are missing some info

OpenStudy (anonymous):

Batter A B G K George 0.319 10.7 46 Jose 0.320 11.4 45 Dale 0.298 12.3 52 Michael 0.337 15.2 48

OpenStudy (anonymous):

A = (b+k)/g A*g = b + k K = A*g - b you can put the numbers in.

OpenStudy (anonymous):

@Deldrickg clear

OpenStudy (anonymous):

@best.shakir Can You Explain A Little More With The Numbers?